ADNAN MUHAMMED

Grade 12,

Find the locus of the middle points of chords of given circle x 2 +y 2 =a 2 which subtends a right angle at the fixed point (p,q).

Find the locus of the middle points of chords of given circle x2+y2=a2 which subtends a right angle at the fixed point (p,q).
 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
locus of the middle points of chords of given circle x2+y2=a2 which subtends a right angle at the fixed point (p,q) is:
T = S1
xx1 + yy1 – a^2 = p^2 + q^2 – a^2
it becomes,
px + qy = p^2 + q^2 therfore.